SEIBERG–WITTEN MONOPOLES ON SEIFERT FIBERED SPACES

TOMASZ S. MROWKA PETER S. OZSVÁTH BAOZHEN YU

Abstract. In this paper, we investigate the Seiberg-Witten gauge theory for Seifert fibered spaces. The monopoles over these three-manifolds, for a particular choice of metric and perturbation, are completely described. Gradient flow lines between monopoles are identified with holomorphic data on an associated ruled surface, and a dimension formula for such flows is calculated.